y = logx base 10 + log10 base x + logx  base x + log10 base 10 find dy/dx

Grade:11

1 Answers

venkat
6 years ago
The answer is simple.
The question can be written as
y=log x/log 10 + log 10/log x +1+1 [Since logxbasex=log x/log x = 1]
On differentiating you get,
dy/dx=1/(xlog 10)  +[-{log 10/(log x)2}*1/x]  +0  +0
dy/dx=1/(xlog 10)  +  [-log 10/x(logx)2]